mang eloy decided to convert his vacant lot into a rectangular garden to earn money for his family. the lot has an area of 84square meter and a perimeter of 25 meter. what would be the length and width of manoy elo'y garden?

Found 2 solutions by ankor@dixie-net.com, greenestamps:
Answer by ankor@dixie-net.com(22740)

.
the lot has an area of 84 square meter and a perimeter of 25 meter. what would be the length and width of manoy elo'y garden?
:
2L + 2w = 25
2L = 25 - 2w
divide equation by 2
L = 12.5 - w
and
L*w = 84
replace L with (12.5-w)
w(12.5-w) = 84
12.5w - w^2 = 84
a quadratic equation
0 = w^2 - 12.5w + 84
Has no real roots, so unsolvable
:
The reality is that you can't get 84 sq meters with that perimeter

Answer by greenestamps(13200)



The given numbers are incompatible; there is no solution.

For a given perimeter, the rectangle with the largest area is a square.

With a perimeter of 25m, the side length of a square would be 6.25m.

The area of a square with side length 6.25m is far less than the required 84 square meters.